OpenStudy (anonymous):

Fan And Medal~~~~~~~ I'm bringing Simple Math Back Simplify the expression by combining like terms. 7p+6(4-9p) I know the answer please show your work

OpenStudy (anonymous):

@danieljoel

OpenStudy (anonymous):

6(4-9p) = (6x4) + ( 6x(-9p)) = ?

OpenStudy (anonymous):

-47p+24

OpenStudy (anonymous):

9x6 = 54 not 57 :D

OpenStudy (anonymous):

so it's -54p + 24

OpenStudy (anonymous):

can you say the answer?

OpenStudy (anonymous):

The answer sheet for my test says -47p+24 But your making a lot of sense

OpenStudy (anonymous):

you know why? because -54 + 7 = -47 :)

OpenStudy (anonymous):

Mathway aloso says -47p+24 Idk how bro

OpenStudy (anonymous):

Oh lol

OpenStudy (anonymous):

so the final answer is -47p + 24 ok?

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Lol Yes I appreciate your Help, Sos101 logging out capt'n

OpenStudy (anonymous):

Yo're welcome :)